### v1

@@ 1. Introduction
With the development of large language models (LLMs) and vision language models (VLMs), research on agents across various domains is booming (Li et al., 2025; Liu et al., 2025; He et al., 2024; Lai et al., 2024; Lu et al., 2024; Sun et al., 2022; You et al., 2024; Chen et al., 2024a; Li et al., 2024; lu2021text2event; deng2018visual), especially those focusing on graphical user interface (GUI) agents in personal computer (PC) and mobile phone scenarios. A number of early works have explored this direction. For example, AppAgent (Zhang et al., 2023) introduces a multimodal agent for smartphone app control using LLMs. Similarly, Mobile-Agent (Wang et al., 2024d) proposes a control framework for mobile apps, and UI-TARS (Qin et al., 2025) employs DPO strategies for model training. AutoGLM (Liu et al., 2024) offers an agent system for autonomous GUI control on web browsers and mobile devices. In addition, CogAgent (Hong et al., 2024), a visual language model, excels at GUI navigation by processing screenshot details. Building on these foundations, powerful models like Qwen-VL-2.5 (Bai et al., 2025; Wang et al., 2024a) enable agents to flexibly manipulate mobile phones, showing great potential in enhancing digital interaction efficiency.
@@ 45 EN 9
Mobile-Eval (Wang et al., 2024d)
@@ 33 EN 10
Mobile-Eval-v2 (Wang et al., 2024c)
@@ 2.1. LLM-based Agent Systems
LLMs and VLMs are significantly advancing Artificial General Intelligence (AGI) through enhanced intelligent agent capabilities (Cai et al., 2025; Chen et al., 2025; Ma et al., 2024; Zeng et al., 2021; Shen et al., 2024; Gao et al., 2024). Frameworks like AutoGPT (Yang et al., 2023) and Metagpt (Hong et al., 2023) pioneer multi-agent collaboration through standardized procedures. In mobile agent systems, AppAgentX (Jiang et al., 2025) enhancing efficiency via memory-driven shortcuts. The Mobile-Agent series (Wang et al., 2024d, c, 2025) progresses from vision-based (OCR/CLIP) approaches to hierarchical multi-agent systems, while VisionTasker (Song et al., 2024) interprets UI screenshots into structured natural language for LLM task planning, DroidBot GPT (Wen et al., 2023) converts GUI states into prompts for autonomous action sequencing. Advanced frameworks like CoCo-Agent (Ma et al., 2024) and CoAT (Zhang et al., 2024b) integrate perception-action mechanisms for precise decision-making, complemented by AutoDroid (Wen et al., 2024a) and AutoDroid-V2 (Wen et al., 2024b)’s hybrid cloud-device architectures for efficient execution. However, most existing frameworks have been conducted in limited scenarios and simple tasks.
@@ 2.2. Benchmarks of Mobile Agents
The development of mobile agent benchmarks has advanced to address increasingly complex GUI interactions (Gu et al., 2025). PIXELHELP (Li et al., 2020) established the foundation by mapping natural language instructions to executable UI actions using task accuracy metrics. ANDROIDLAB (Xu et al., 2024) formalized Android agent evaluation with XML and SoM operational modes, covering 138 tasks across nine applications. Mobile-Bench (Deng et al., 2024) extended evaluation to multi-app workflows, combining API and UI operations with CheckPoint metrics for planning analysis. MobileSafetyBench (Lee et al., 2024) incorporated layered risk assessments including legal compliance tests to address safety concerns. SPA-BENCH (Chen et al., 2024b) expanded evaluations to multilingual and cross-app scenarios with efficiency metrics. MobileAgentBench (Wang et al., 2024b) provided a fully autonomous evaluation process on real Android devices. LlamaTouch (Zhang et al., 2024a) enhanced evaluation precision through state-aware annotations. GTArena (Zhao et al., 2024) developed GUI defect detection using Transition Tuples. Mobile-Eval-E (Wang et al., 2025) specifically targets reasoning-intensive multi-app workflows.
@@ 4.1. Model Training
Aider is a lightweight, plug-and-play module, fine-tuned model based on Qwen2.5-VL-3B (Bai et al., 2025). Firstly, we use GPT-4 with specific prompts to generate the required instructions, such as unethical or vague ones. We then hire five expert taggers to manually capture mobile screenshots and verify the consistency of task classification. This process resulted in 1,000 high-quality question-and-answer pairs, including responses to unethical instructions. This dataset was subsequently utilized for model training to fine-tune the full parameters in one epoch, thereby developing our Aider model
@@ 5.1. Implementation Details
We evaluate a spectrum of influential open-source mobile agent frameworks, including Mobile-Agent (Wang et al., 2024d), Mobile-Agent-V2 (Wang et al., 2024c), and Mobile-Agent-E (Wang et al., 2025), spanning both single-agent and multi-agent architectures to comprehensively represent cutting-edge advancements in current mobile agents frameworks.
@@
We evaluate six open-source and closed-source large vision language models as backbones for the three agent frameworks, including GPT-4o [^3] https://openai.com/index/, Gemini-2.0-pro [^4] https://deepmind.google/technologies/gemini/flash/, Claude-3.5-Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet and Qwen2.5-VL series (72B,7B,3B) (Bai et al., 2025).
@@ C.2. Model Training
We fine-tune the Qwen2.5-VL-3B-Instruct model using LoRA. Training is conducted for 3 epochs on 1000 samples with a maximum sequence length of 2048. We use the AdamW optimizer with a cosine scheduler, a learning rate of $5\times 10^{-5}$, and an effective batch size of 16 (2 per device, 8 gradient accumulation steps). Model training prompt is shonw in Table 8
@@ Appendix D Visualization of MVISU-Bench
We further analyze the distribution of Chinese and English apps in MVISU-Bench, as well as the frequency of instructions within the dataset. The distributions of Chinese and English apps are shown in Figure 7 and Figure 8, respectively. The instruction word cloud statistics are presented in Figure 9 and Figure 10. The full benchmark dataset of English and Chinese instructions are provided in Table LABEL:tab:cn and Table LABEL:tab:en.

### v2

@@ 1. Introduction
With the development of large language models (LLMs) and vision language models (VLMs), research on agents across various domains is booming (Li et al., 2025; Liu et al., 2025; He et al., 2024; Lai et al., 2024; Lu et al., 2024; Sun et al., 2022; You et al., 2024; Chen et al., 2024a; Li et al., 2024; Lu et al., 2021; Deng et al., 2018), especially those focusing on graphical user interface (GUI) agents in personal computer (PC) and mobile phone scenarios. A number of early works have explored this direction. For example, AppAgent (Zhang et al., 2023) introduces a multimodal agent for smartphone app control using LLMs. Similarly, Mobile-Agent (Wang et al., 2024c) proposes a control framework for mobile apps, and UI-TARS (Qin et al., 2025) employs DPO strategies for model training. AutoGLM (Liu et al., 2024) offers an agent system for autonomous GUI control on web browsers and mobile devices. In addition, CogAgent (Hong et al., 2024), a visual language model, excels at GUI navigation by processing screenshot details. Building on these foundations, powerful models like Qwen-VL-2.5 (Team, 2025, 2024) enable agents to flexibly manipulate mobile phones, showing great potential in enhancing digital interaction efficiency.
@@ 45 EN 9
Mobile-Eval (Wang et al., 2024c)
@@ 33 EN 10
Mobile-Eval-v2 (Wang et al., 2024b)
@@ 2.1. LLM-based Agent Systems
LLMs and VLMs are significantly advancing Artificial General Intelligence (AGI) through enhanced intelligent agent capabilities (Cai et al., 2025; Chen et al., 2025; Ma et al., 2024; Zeng et al., 2021; Shen et al., 2024; Gao et al., 2024). Frameworks like AutoGPT (Yang et al., 2023) and Metagpt (Hong et al., 2023) pioneer multi-agent collaboration through standardized procedures. In mobile agent systems, AppAgentX (Jiang et al., 2025) enhancing efficiency via memory-driven shortcuts. The Mobile-Agent series (Wang et al., 2024c, b, 2025) progresses from vision-based (OCR/CLIP) approaches to hierarchical multi-agent systems, while VisionTasker (Song et al., 2024) interprets UI screenshots into structured natural language for LLM task planning, DroidBot GPT (Wen et al., 2023) converts GUI states into prompts for autonomous action sequencing. Advanced frameworks like CoCo-Agent (Ma et al., 2024) and CoAT (Zhang et al., 2024b) integrate perception-action mechanisms for precise decision-making, complemented by AutoDroid (Wen et al., 2024a) and AutoDroid-V2 (Wen et al., 2024b)’s hybrid cloud-device architectures for efficient execution. However, most existing frameworks have been conducted in limited scenarios and simple tasks.
@@ 2.2. Benchmarks of Mobile Agents
The development of mobile agent benchmarks has advanced to address increasingly complex GUI interactions (Gu et al., 2025). PIXELHELP (Li et al., 2020) established the foundation by mapping natural language instructions to executable UI actions using task accuracy metrics. ANDROIDLAB (Xu et al., 2024) formalized Android agent evaluation with XML and SoM operational modes, covering 138 tasks across nine applications. Mobile-Bench (Deng et al., 2024) extended evaluation to multi-app workflows, combining API and UI operations with CheckPoint metrics for planning analysis. MobileSafetyBench (Lee et al., 2024) incorporated layered risk assessments including legal compliance tests to address safety concerns. SPA-BENCH (Chen et al., 2024b) expanded evaluations to multilingual and cross-app scenarios with efficiency metrics. MobileAgentBench (Wang et al., 2024a) provided a fully autonomous evaluation process on real Android devices. LlamaTouch (Zhang et al., 2024a) enhanced evaluation precision through state-aware annotations. GTArena (Zhao et al., 2024) developed GUI defect detection using Transition Tuples. Mobile-Eval-E (Wang et al., 2025) specifically targets reasoning-intensive multi-app workflows.
@@ 4.1. Model Training
Aider is a lightweight, plug-and-play module, fine-tuned model based on Qwen2.5-VL-3B (Team, 2025). Firstly, we use GPT-4 with specific prompts to generate the required instructions, such as unethical or vague ones. We then hire five expert taggers to manually capture mobile screenshots and verify the consistency of task classification. This process resulted in 1,000 high-quality question-and-answer pairs, including responses to unethical instructions. This dataset was subsequently utilized for model training to fine-tune the full parameters in one epoch, thereby developing our Aider model
@@ 5.1. Implementation Details
We evaluate a spectrum of influential open-source mobile agent frameworks, including Mobile-Agent (Wang et al., 2024c), Mobile-Agent-V2 (Wang et al., 2024b), and Mobile-Agent-E (Wang et al., 2025), spanning both single-agent and multi-agent architectures to comprehensively represent cutting-edge advancements in current mobile agents frameworks.
@@
We evaluate six open-source and closed-source large vision language models as backbones for the three agent frameworks, including GPT-4o [^3] https://openai.com/index/, Gemini-2.0-pro [^4] https://deepmind.google/technologies/gemini/flash/, Claude-3.5-Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet and Qwen2.5-VL series (72B,7B,3B) (Team, 2025).
@@ C.2. Model Training
We fine-tune the Qwen2.5-VL-3B-Instruct model using LoRA. Training is conducted for 3 epochs on 1000 samples with a maximum sequence length of 2048. We use the AdamW optimizer with a cosine scheduler, a learning rate of $5\times 10^{-5}$, and an effective batch size of 16 (2 per device, 8 gradient accumulation steps). Model training prompt is shonw in Table 8.
@@ Appendix D Visualization of MVISU-Bench
We further analyze the distribution of Chinese and English apps in MVISU-Bench, as well as the frequency of instructions within the dataset. The distributions of Chinese and English apps are shown in Figure 7 and Figure 8, respectively. The instruction word cloud statistics are presented in Figure 9 and Figure 10.